# Adaptive Deadline and Batch Layered Synchronized Federated Learning

### v1

@@ 3.4 Discussion
The cost function in Problem 2 reflects a fundamental trade-off between the terms $B$ and $C_{t}$, which capture distinct sources of error. The term $B=\frac{1}{mU^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{P_{u}}+6\rho_{s}\Gamma$ corresponds to the variance induced by stochastic gradients, and is inversely proportional to the global batch scaling factor $m$. Thus, increasing $m$ (i.e., using larger batches) reduces $B$, reducing the user’s SGD variance. However, larger batches require more computation per user, potentially reducing the number of clients able to contribute full updates within a fixed deadline. This effect is captured by $C_{t}$, which quantifies the variance due to partial participation and depends on the regularized incomplete gamma function. As $m$ increases, $C_{t}$ worsens due to fewer layers being computed per user under time constraints as in Requirement R2. In addition, the per-round deadline $T_{t}^{d}$ must be judiciously allocated across rounds to align with the learning rate schedule and batch size configuration, ensuring that each round maximally leverages the available computational budget while preserving convergence efficiency.
A natural extension of ADEL-FL is to also optimize the number of rounds $R$, in addition to $T_{t}^{d}$ and $m$. In practical deployments, both the number of rounds and the overall latency are often constrained. Optimizing over $R$ introduces an additional degree of flexibility, allowing the system to determine the optimal number of global rounds given the available time budget. This joint optimization problem—-over the number of rounds $R$, per-round deadlines $\{T_{t}^{d}\}$, and batch scaling $m$—-could be formulated as a mixed-integer constrained program or tackled with adaptive scheduling heuristics, replacing Problem 2 in line 2 of the algorithm. Incorporating this optimization into ADEL-FL would further enhance its applicability to time-critical FL scenarios.
4 Experimental Results
[…]
@@ 4 Experimental Results
Figure 2: Deadline allocation and convergence curves for MNIST using an inverse decaying learning rate. $(a)$ Adaptive deadline per round for ADEL-FL (MLP). $(b)$ Convergence (validation accuracy vs. training time), MLP. $(c)$ Convergence, CNN.
MNIST
@@
We first evaluate on the MNIST dataset [12] (handwritten $28\times 28$ grayscale digits). The data are partitioned uniformly at random across $U=20$ users (each client holds $1/20$ of the training set of 60,000 images). We choose the total training time $T_{\max}$ and number of rounds $R$ such that, under a standard local minibatch size, the average fraction of each model updated per round (i.e. the average backpropagation depth) is $50\%$ of the network layers. We consider two model architectures of different capacity: (1) MLP: a simple fully-connected network with two hidden layers (32 and 16 neurons) and a softmax output; (2) CNN: a small CNN with two $5\times 5$ convolutional layers (each followed by max-pooling and ReLU [29]) and two dense layers, ending in a softmax output.
@@
Next, we evaluate on the CIFAR-10 simulating an FL system with $U=30$ users, each receiving an equal partition of the 50,000 training images. In these experiments, we focus on deeper neural network models. We train two architectures of the VGG family [30]: VGG11 and VGG13, which have 8 and 10 convolutional layers (respectively) followed by 3 fully-connected layers. These networks are substantially deeper than the MNIST models, providing a more challenging and heterogeneous workload. We set the global time budget $T_{\max}$ and rounds $R$ such that the average local computation per round reaches about $85\%$ of each model’s full depth for a baseline batch size – i.e., on average clients nearly complete a full pass through the network each round under the time limit. Results are shown in Figure 3. ADEL-FL consistently outperforms SALF, Drop-Stragglers, and Wait Stragglers, achieving faster convergence while maintaining a higher final accuracy (69% for ADEL-FL  vs 65% for SALF). As in MNIST, the deadline allocation decreases over time alongside the decaying learning rate, efficiently balancing training speed and convergence.
@@
Figure 3: Deadline allocation and convergence curves for CIFAR10 using an inverse decaying learning rate. $(a)$ Adaptive deadline per round for ADEL-FL (VGG11). $(b)$ Convergence, VGG11. $(c)$ Convergence, VGG13.
@@
Figure 4: Robustness results, CIFAR-10 VGG11. $(a)$ Convergence with $\ell_{2}$ regularization. $(b)$ Convergence with constant learning rate.
Robustness to Learning Assumptions
To validate ADEL-FL under conditions where optimization assumptions are violated, we perform two studies on CIFAR-10 VGG11, reported in Figure 4. First, we introduce $\ell_{2}$ regularization on local objectives. Second, we fix the learning rate $\eta_{t}=\eta_{0}$ instead of using a decaying profile. These modifications deviate from theoretical conditions assumed for convergence in Theorem 1. Nevertheless, ADEL-FL maintains advantages over SALF, Wait-Stragglers and Drop-Stragglers. The convergence remains stable and outperforms the other existing solutions compared, demonstrating that ADEL-FL’s design is robust even when the actual optimization dynamics differs moderately from the assumed model.
@@
Table 1: Final test accuracy (%) of ADEL-FL and baseline methods on CIFAR-10 (VGG11) under varying total training time budgets $T_{\max}$.
Training Time
Table 1 compares final accuracy across different total training time budgets. Even under stringent time limits, ADEL-FL preserves superior convergence, highlighting its efficient use of limited computation. As the total training time $T_{\max}$ increases, all methods generally benefit from improved model accuracy due to deeper local computations. However, ADEL-FL consistently maintains a noticeable gap over SALF, Wait-Stragglers and Drop-Stragglers across all budgets. In the low-budget regime, ADEL-FL is particularly advantageous, achieving much better performance while effectively handling the computation resources. In higher budget regimes, although the gap slightly narrows, ADEL-FL continues to converge faster and achieves superior final performance.
5 Conclusion
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Appendix A - Proof of Lemma 1
We start by defining $d_{t}^{u}$ as the depth reached by the user $u$ in the round $t$, up to the deadline $T_{t}^{d}$. When Assumption A4 holds, backpropagation can be modeled as sequential layer gradient computation with exponential service time. We get a Poisson process, where $d_{t}^{u}$ is closely related to the number of layers processed:
$d_{t}^{u}=L+1-\min\{z_{t}^{u},L\},\quad\text{where}\quad z_{t}^{u}\sim\text{Poiss}\bigg{(}\frac{T_{t}^{d}}{m}\bigg{)}.$
Recall that $p_{l}\triangleq P(|\mathcal{U}_{t}^{l}|=0)$ denotes the probability that no user contributes an update for layer $l$ at round $t$. We use the i.i.d (by $u$) property of Assumption A4 to write:
$P(|\mathcal{U}_{t}^{l}|=0)=\bigcap_{u=1}^{U}P(z_{t}^{u}\leq L-l)=\prod_{u=1}^{U}P(z_{t}^{u}\leq L-l)=\left(P(z_{t}^{1}\leq L-l)\right)^{U}.$
The probability mass function (PMF) of $X\sim\text{Poiss}(\lambda)$ is given by:
@@ A.1 Appendix A - Proof of Lemma 1
We denote $\lambda_{t}\triangleq\frac{T_{t}^{d}}{m}$ and get:

### v2

@@ III-D Discussion
The ADEL-FL framework aims to minimize convergence error by adaptively managing the interplay between local workload and global timing constraints. This is achieved by optimizing the trade-off between stochastic gradient noise and partial participation variance, as formalized in the cost function of Problem 2. The cost function reflects a fundamental trade-off between the terms $B_{t}$ and $C_{t}$, which capture distinct sources of error. The term $B_{t}$ corresponds to the variance induced by stochastic gradients, and is a monotonically decreasing function of the global batch scaling factor $m$. Thus, increasing $m$ (i.e., using larger batches) reduces $B_{t}$, reducing the user’s SGD variance. However, larger batches require more computation per user, potentially reducing the number of clients able to contribute full updates within a fixed deadline. This effect is captured by $C_{t}$, which quantifies the variance due to partial participation and depends on the regularized incomplete gamma function. As $m$ increases, $C_{t}$ worsens due to fewer layers being computed per user under time constraints as in Requirement R2. In addition, the per-round deadline $T_{t}^{\rm d}$ must be judiciously allocated across rounds to align with the learning rate schedule and batch size configuration, ensuring that each round maximally leverages the available computational budget while preserving convergence efficiency. A natural extension of ADEL-FL is to also optimize the number of rounds $R$, in addition to $T_{t}^{\rm d}$ and $m$. In practical deployments, both the number of rounds and the overall latency are often constrained. Optimizing over $R$ introduces an additional degree of flexibility, allowing the system to determine the optimal number of global rounds given the available time budget. This joint optimization problem (over the number of rounds $R$, per-round deadlines $\{T_{t}^{\rm d}\}$, and batch scaling $m$) could be formulated as a mixed-integer constrained program or tackled with adaptive scheduling heuristics, replacing Problem 2 in Step 2 of the algorithm. Incorporating this optimization into ADEL-FL would further enhance its applicability to time-critical FL scenarios. Furthermore, while this work assumes a deterministic communication model for tractability, incorporating the extension to stochastic or time-varying communication models remains an important direction for future research.
@@
Figure 2: Deadline allocation and convergence curves for MNIST using an inverse decaying learning rate. $(a)$ Adaptive deadline per round for ADEL-FL (MLP).(b)(b) Convergence, MLP. $(c)$ Adaptive deadline per round for ADEL-FL (CNN).
IV Experimental Study
We empirically evaluate ADEL-FL for training image classification models using MNIST [33] and CIFAR-10 [34]. We consider a series of experiments of increasing complexity. In each case, we compare ADEL-FL against four baseline schemes: $(i)$ Wait Stragglers: the standard federated averaging with full client participation (synchronous FedAvg with the server waiting for all devices each round) [1]; $(ii)$ Drop-Stragglers: synchronous FedAvg with a fixed per-round deadline that drops any late client updates [17]; $(iii)$ SALF: the straggler-aware FL method suggested in  [31], which aggregates partial gradients from slow devices but uses a fixed deadline and batch size; and $(iv)$ HeteroFL: which trains local submodels with reduced width according to each client’s capability, allowing efficient aggregation into a full global model [30]. All methods are evaluated under a fixed total training time budget $T_{\max}$ and a fixed number of communication rounds $R$ (per the constraints in Requirements R1R2). Experiments are conducted on a Linux server with NVIDIA A30 GPU. To emulate heterogeneous device speeds in a controlled manner, we sample each client’s computation time per layer from an exponential distribution (as in Model Formulation B1). We begin with image classification tasks on MNIST and CIFAR-10 using an inverse-decay learning rate schedule $\eta_{t}=\eta_{0}/(1+t)$ with initial $\eta_{0}\in\{0.05,\,0.1,\,0.5,\,1.0\}$ to assess baseline convergence behavior. Next, we analyze stability and optimization dynamics under constant learning rate and adding $\ell_{2}$ regularization, deviating from the assumptions of Theorem 1. Finally, we present a comparative table summarizing performance across varying total training times, highlighting ADEL-FL’s efficiency under strict time constraints.
IV-A MNIST
@@
We first evaluate on the MNIST dataset [33] (handwritten $28\times 28$ grayscale digits). We set the total training time $T_{\max}$ and number of rounds $R$ such that, under a standard local mini-batch size, the average fraction of each model updated per round (i.e. the average backpropagation depth) is $50\%$ of the network layers. We consider two model architectures of different capacity: $(1)$ MLP: a simple fully-connected network with two hidden layers (32 and 16 neurons) and a softmax output; and $(2)$ CNN: a small CNN with two $5\times 5$ convolutional layers (each followed by max-pooling and ReLU) and two dense layers, ending in a softmax output.
@@ IV-B CIFAR-10
Next, we evaluate on CIFAR-10, simulating an FL system with $U=30$ users. The data are partitioned across the users using a Dirichlet distribution with concentration parameter $\alpha=0.5$, following the standard non-IID construction of [49]. This procedure samples client-specific label proportions from a Dirichlet prior, producing heterogeneous yet controllable data heterogeneity across users. Consequently, each client receives a distinct subset of the 60,000 training images that reflects a moderately non-IID allocation of the global dataset. In these experiments, we focus on deeper neural network models. We train two architectures of the VGG family [50]: VGG11 and VGG13, which have 8 and 10 convolutional layers (respectively) followed by 3 fully-connected layers. These networks are substantially deeper than the MNIST models, providing a more challenging and heterogeneous workload. We set the global time budget $T_{\max}$ and rounds $R$ such that the average local computation per round reaches about $85\%$ of each model’s full depth for a baseline batch size, i.e., on average clients nearly complete a full pass through the network each round under the time limit. The resulting allocations and validation accuracies are reported in Figure 3. We observe that ADEL-FL consistently outperforms SALF, Drop-Stragglers, and Wait Stragglers, achieving faster convergence while maintaining a higher final accuracy (66% for ADEL-FL vs 53% for SALF). As in MNIST, the deadline allocation decreases over time alongside the decaying learning rate, efficiently balancing training speed and convergence.
@@
Figure 3: Deadline allocation and convergence curves for CIFAR-10 using an inverse decaying learning rate. $(a)$ Adaptive deadline per round for ADEL-FL (VGG11).(b)(b) Convergence, VGG11. $(c)$ Adaptive deadline per round for ADEL-FL (VGG11).(d)(d) Convergence, VGG11.
IV-C Robustness to Learning Assumptions
To validate ADEL-FL under conditions where optimization assumptions are violated, we perform four studies on CIFAR-10 VGG11, reported in Figure 4. First, we introduce $\ell_{2}$ regularization on local objectives. Using $\ell_{2}$ regularization has been shown to improve generalization and to stabilize training in modern deep networks, including when adaptive optimizers are used. Empirical studies demonstrate that decoupled weight decay improves final accuracy and yields more reliable behavior across datasets and architectures, which makes it a practical tool to mitigate overfitting on clients with limited local data [51, 1, 29]. Second, we fix the learning rate $\eta_{t}=\eta_{0}$ instead of using a decaying profile. As discussed in [52], maintaining a constant learning rate during early optimization helps SGD preserve gradient noise, enabling broader exploration of the loss landscape and reducing the tendency to converge prematurely to sharp or suboptimal minima.
@@ IV-C Robustness to Learning Assumptions
Figure 4: Robustness results, CIFAR-10 VGG11. $(a)$ Convergence with $\ell_{2}$ regularization. $(b)$ Convergence with constant learning rate. $(c)$ Convergence with $E=3$ local iterations. $(d)$ Convergence with $E=5$ local iterations.
So far, we have focused our formulation and experiments on settings where the users compute a single stochastic gradient over the selected mini-batch, i.e., using the local SGD terminology as in [5], by letting $E$ be the number of local SGD iterations, we so far used $E=1$. In the third and fourth experiments, we vary the number of local SGD iterations before each global update to $E=3$ and $E=5$ respectively. Increasing the number of local iterations $E$ is a logical and widely used strategy in FL because it reduces communication frequency, lowers network load, and allows each global aggregation step to amortize more local computation. Prior work has shown that performing multiple local updates per round significantly decreases communication requirements in practice [1], and that local SGD can maintain strong convergence behavior even under aggressive communication reduction [5]. At the same time, larger $E$ increases the risk of client drift in heterogeneous data regimes [53]. These modifications deviate from theoretical conditions assumed for convergence in Theorem 1. Despite the fact that the considered setting no longer aligns with the assumptions based on which our objective is derived, we observe in Figure 4 that ADEL-FL maintains advantages over SALF, Wait-Stragglers and Drop-Stragglers. Across all stress tests, the method exhibits stable descent behavior and improved final accuracy, suggesting that its synchronization and deadline-adaptive aggregation mechanism compensates for moderate deviations from canonical optimization assumptions. The convergence remains stable and outperforms the other existing solutions compared, demonstrating that ADEL-FL’s design is robust even when the actual optimization dynamics differs moderately from the assumed model. This robustness further indicates that ADEL-FL can be deployed in realistic federated environments where heterogeneous client behaviors often violate idealized assumptions, without sacrificing stability or efficiency.
IV-D Training Time
We conclude our experimental study by evaluating the performance achieved under predefined overall deadlines. Table II compares final accuracy across different total training time budgets for the VGG11 architecture on the CIFAR-10 dataset, using an IID data partition across clients. Even under stringent time limits, ADEL-FL preserves superior convergence, highlighting its efficient use of limited computation. As the total training time $T_{\max}$ increases, all methods generally benefit from improved model accuracy due to deeper local computations. However, ADEL-FL consistently maintains a noticeable gap over SALF, Wait-Stragglers and Drop-Stragglers across all budgets. In the low-budget regime, ADEL-FL is particularly advantageous, achieving much better performance while effectively handling the computation resources. In higher budget regimes, although the gap slightly narrows, ADEL-FL continues to converge faster and achieves superior final performance.
TABLE II: Final test accuracy (%) of ADEL-FL and baseline methods on CIFAR-10 (VGG11) under varying total training time budgets $T_{\max}$.
@@ IV-D Training Time
V Conclusion
@@ Appendix A Technical Appendices and Supplementary Material
A-A Appendix A - Proof of Lemma 1
We initiate the proof by defining $d_{t}^{u}$ as the index of the final layer reached during the backpropagation process by user $u$ in round $t$. Given a model with total depth $L$, this index is determined by the relationship:
$d_{t}^{u}=\max\{L+1-z_{t}^{u},1\},$ (A.1)
where $z_{t}^{u}$ represents the number of layers that would be processed within the effective deadline $T_{t}^{\rm d}-B_{u}$ assuming an unconstrained (potentially infinite-depth) sequence of layers. This formulation is used since $z_{t}^{u}$ is stochastic and unbounded, while the actual model depth is limited to $L$. Under Model Formulation B1, backpropagation can be modeled as sequential layer gradient computation with exponential service time. By the fundamental relationship between the exponential distribution and the Poisson process [54], the number of layers processed $z_{t}^{u}$ in the fixed interval $0,T_{t}^{\rm d}-B_{u}]$ follows a Poisson distribution:
$z_{t}^{u}\sim\text{Poiss}\left(\lambda_{t}^{u}\right),\quad\text{where }\lambda_{t}^{u}=\frac{P_{u}}{S_{t}^{u}}(T_{t}^{\rm d}-B_{u}).$ (A.2)
By substituting $S_{t}^{u}=\Bigg\lfloor mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)\Bigg\rfloor$ from Model Formulation B3 we get:
$\lambda_{t}^{u}=\frac{mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)}{\Bigg\lfloor mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)\Bigg\rfloor}\frac{T_{t}^{\rm d}}{m}$
Recall that $p_{t}^{l}\triangleq P(|\mathcal{U}_{t}^{l}|=0)$ denotes the probability that no user contributes an update for layer $l$ at round $t$. We use the i.i.d (by $u$) property of Model Formulation B1 to write:
$P(|\mathcal{U}_{t}^{l}|=0)$ $=\bigcap_{u=1}^{U}P(z_{t}^{u}\leq L-l)$
$=\prod_{u=1}^{U}P(z_{t}^{u}\leq L-l)=\left(P(z_{t}^{1}\leq L-l)\right)^{U}$
To simplify the bound, we introduce an auxiliary random variable $\tilde{z}_{t}$:
$\tilde{z}_{t}\sim\text{Poiss}\left(\tilde{\lambda}_{t}\right),\quad\text{where }\tilde{\lambda}_{t}=\frac{T_{t}^{\rm d}}{m},$ (A.3)
which provides an upper bound on the zero-contribution probability:
$P(|\mathcal{U}_{t}^{l}|=0)=\left(P(z_{t}^{1}\leq L-l)\right)^{U}\leq\left(P(\tilde{z}_{t}\leq L-l)\right)^{U}.$ (A.4)
This inequality holds because $\lambda_{t}^{u}\geq\tilde{\lambda}_{t}$, and since the Poisson cumulative distribution function is a monotonically decreasing function of its rate, the smaller rate $\tilde{\lambda}_{t}$ provides a conservative upper bound for the zero-contribution probability. The probability mass function (PMF) of $X\sim\text{Poiss}(\lambda)$ is given by:
@@ A-A Appendix A - Proof of Lemma 1
We get:
$P(|\mathcal{U}_{t}^{l}|=0)\leq\left(P(\tilde{z}_{t}\leq L-l)\right)^{U}=\left(\sum_{k=0}^{L-l}\frac{\tilde{\lambda}_{t}^{k}e^{-\tilde{\lambda}_{t}}}{k!}\right)^{U}.$ (A.5)
It can be observed that the summation term in Eq. (A.5) coincides with the regularized upper incomplete gamma function evaluated at a specific argument. The following auxiliary lemma formalizes this identity, with the proof deferred to Section E.
Auxiliary Lemma. For an integer $s\geq 1$,
$Q(s,x)=\sum_{k=0}^{s-1}\frac{x^{k}e^{-x}}{k!},$ (A.6)
where $Q(s,x)\triangleq\frac{1}{\Gamma(s)}\int_{x}^{\infty}t^{s-1}e^{-t}dt$ is the regularized upper incomplete gamma function.
Using the auxiliary lemma with s1=Lls-1=L-l and $x=\frac{T_{t}^{\rm{d}}}{m}$, Eq. (A.5) can be written as
$p_{t}^{l}=P(|\mathcal{U}_{t}^{l}|=0)\leq Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U},$
which completes the proof of Lemma 1. $\square$
A-B Appendix B - Proof of Lemma 2
$\mathbb{E}\Big[\tilde{{\bm{w}}}_{t+1}^{l}\Big]$ $=\mathbb{E}\Big[\mathbb{E}\Big[\tilde{{\bm{w}}}_{t+1}^{l}\Big||\mathcal{U}_{t}^{l}|\Big]\Big]$
$=\sum_{K=0}^{U}P(|\mathcal{U}_{t}^{l}|=K)\mathbb{E}\Big[\tilde{{\bm{w}}}_{t+1}^{l}\Big||\mathcal{U}_{t}^{l}|=K\Big].$
Substituting the probability of $|\mathcal{U}_{t}^{l}|=0$, denoted as $p_{t}^{l}$, we get:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]$ $=p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}+\sum_{K=1}^{U}P(|\mathcal{U}_{t}^{l}|=K)\frac{1}{1-p_{t}^{l}}\cdot$
$\left(\frac{1}{K}\mathbb{E}\left[\sum_{u\in\mathcal{U}_{t}^{l}}{\bm{w}}_{u,t}^{l}\middle||\mathcal{U}_{t}^{l}|=K\right]-p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}\right).$
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]=p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}+\frac{{\bm{w}}_{t+1}^{l}-p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}}{1-p_{t}^{l}}\sum_{K=1}^{U}P(|\mathcal{U}_{t}^{l}|=K).$
Using:
$\sum_{K=1}^{U}P(|\mathcal{U}_{t}^{l}|=K)=1-P(|\mathcal{U}_{t}^{l}|=0)=1-p_{t}^{l},$
we get:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}^{l}]=p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}+{\bm{w}}_{t+1}^{l}-p_{t}^{l}\tilde{{\bm{w}}}_{t}^{l}={\bm{w}}_{t+1}^{l}.$
Since this equality holds for every layer $l\in\{1,\dots,L\}$, and the global model $\tilde{{\bm{w}}}_{t+1}$ is formed by stacking the layer-wise vectors $\tilde{{\bm{w}}}^{l}_{t+1}$, it follows that the full model update is also unbiased:
$\mathbb{E}[\tilde{{\bm{w}}}_{t+1}]={\bm{w}}_{t+1},$
completing the proof. $\square$
A-C Appendix C - Proof of Lemma 3
$v_{t+1}^{l}$ $\triangleq\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]$ (A.8)
$=P\Big[|\mathcal{U}_{t}^{l}|=0\Big]\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big||{\mathcal{U}}_{t}^{l}|=0\Big]$
For the first summand, it holds that:
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big||{\mathcal{U}}_{t}^{l}|=0\Big]=\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]$ (A.9)
$=$ $\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}+{\bm{w}}_{t}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\Big]$
$=$ $\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big]+\mathbb{E}\Big[\|{\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big]$
$\quad+2\mathbb{E}\Big[\big(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}\big)^{T}\big({\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}\big)\Big]$
$\stackrel{{\scriptstyle(a)}}{{=}}$ $v_{t}^{l}+\mathbb{E}\Big[\|{\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $v_{t}^{l}+\sum_{u=1}^{U}\frac{1}{U}\mathbb{E}\Big[\|{\bm{w}}_{u,t}^{l}-{\bm{w}}_{t}^{l}\|^{2}\Big]$
$=$ $v_{t}^{l}+\sum_{u=1}^{U}\frac{1}{U}\mathbb{E}\Big[\|\eta_{t}\nabla F_{u}({\bm{w}}_{t};i_{t}^{u})\|^{2}\Big]\stackrel{{\scriptstyle(c)}}{{\leq}}\;v_{t}^{l}+\eta_{t}^{2}G^{2}.$
where (a) follows by conditioning on the current batches $\{i_{t}^{u}\}_{u=1}^{U}$. Given $\{i_{t}^{u}\}_{u=1}^{U}$, the quantity ${\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}$ is deterministic, and since $\tilde{{\bm{w}}}_{t}^{l}$ is determined by the aggregation randomness up to round t1t-1 and is independent of the current batches, Lemma 2 implies $\mathbb{E}\big[\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}\,\big|\,\{i_{t}^{u}\}_{u=1}^{U}\big]=\mathbf{0}$, which yields $\mathbb{E}\Big[\big(\tilde{{\bm{w}}}_{t}^{l}-{\bm{w}}_{t}^{l}\big)^{T}\big({\bm{w}}_{t+1}^{l}-{\bm{w}}_{t}^{l}\big)\Big]=0$. Step (b) follows from the convexity of $\|\cdot\|^{2}$, and (c) follows from Assumption A3.
For the second summand:
$\mathbb{E}\Big[\|\tilde{{\bm{w}}}_{t+1}^{l}-{\bm{w}}_{t+1}^{l}\|^{2}\ \Big||{\mathcal{U}}_{t}^{l}|=K\Big]$